### v1

# DualFluidNet: an Attention-based Dual-pipeline Network for Accurate and Generalizable Fluid-solid Coupled Simulation

Abstract
Fluid motion can be considered as point cloud transformation when adopted by a Lagrangian description. Compared to traditional numerical analysis methods, using machine learning techniques to learn physics simulations can achieve near accuracy, while significantly increasing efficiency. In this paper, we propose an innovative approach for 3D fluid simulations utilizing an Attention-based Dual-pipeline Network, which employs a dual-pipeline architecture, seamlessly integrated with an Attention-based Feature Fusion Module. Unlike previous single-pipeline approaches, we find that a well-designed dual-pipeline approach achieves a better balance between global fluid control and physical law constraints. Furthermore, we design a Type-aware Input Module to adaptively recognize particles of different types and perform feature fusion afterward, such that fluid-solid coupling issues can be better dealt with. The experiments show that our approach significantly increases the accuracy of fluid simulation predictions and enhances generalizability to previously unseen scenarios. We demonstrate its superior performance over the state-of-the-art approaches across various metrics.
1 Introduction
Figure 1: Armadillo in the flowing water, with 138k particles.
Comprehending physics is essential for enhancing our understanding of the environment and our interactions with it. Traditional fluid simulations have limitations in practical applications due to their high computational demands. Recent studies [29, 3, 22, 17, 26] have highlighted the potential of physics-based neural networks as a promising approach for comprehending intricate natural phenomena from data. Inspired by Smoothed Particle Hydrodynamics (SPH) methods [6, 11], which represent fluids as smoothed particles, several networks [13, 19, 24, 27, 8] have been developed for particle-based simulations. The prediction of physical properties using neural networks is always treated as a regression problem, enabling the effective learning and approximation of physical phenomena.
While previous methods have demonstrated excellent performance in fluid simulation, we have observed that single-path networks always struggle to maintain a balance between overall particle control accuracy and adherence to physical laws. The dilemma in single-path methods turns into a trade-off issue. To deal with this problem, we propose a Dual-Pipeline Network architecture, consisting of a main pathway and a constraint-guided pathway. The main pathway is responsible for establishing a global context and stabilizing the fluid. The constraint-guided pathway incorporates enforced constraints, learning fluid motion patterns under the physical laws. Our network seamlessly fuses the two learning ways through a meticulously designed approach, achieving a harmonious balance between these two pathways and unlocking their maximum potential. Instead of manually determining the proportion of importance for the two pathways, we employ an innovative Attention-based Feature Fusion Module, allowing the network to autonomously learn the optimal fusion between the two paths.
Although previous networks appear to perform well in handling fluid point clouds, there hasn’t been an effective method for handling the mixed input of fluid and solid. Common practices involve simply concatenating the two inputs and feeding them into the network [27], or passing each through a few convolutional layers before concatenation [19]. These straightforward input concatenations are insufficient to enable the network to effectively distinguish between the features of fluid and solid particles. Nevertheless, the fluid-solid coupling is crucial for enhancing fluid simulation performance. Common feature fusion methods are frequently utilized for multimodal and multiscale fusion problems [31, 15, 5]. In the context of fluid simulation, our input comprises two entirely different features: liquid particles and solid particles. Additionally, the particle count varies in each input. We propose an innovative approach to facilitate better differentiation between fluid and solid particles in the input, enabling the network to more accurately and reliably compute their interactions during collisions. We verified that the Type-aware Input Module can be easily transferred to other network models and plays a crucial role in enhancing fluid simulation performance.
We conducted experiments on multiple 3D liquid datasets, and the results demonstrated that our network outperformed existing methods across various performance metrics, surpassing even the state-of-the-art network [19]. Furthermore, we validated the strong generalization of our network model in more challenging scenarios.
2 Related work
Comprehending physics is crucial for making sense of our surroundings and interacting with them more effectively. Smoothed Particle Hydrodynamic (SPH), representing fluids as smoothed particles, is a meshless Lagrangian method commonly employed for simulating fluid dynamics. This perspective inherently conserves fluid mass since the particle count remains constant and efficiently focuses computational resources on the fluid’s location. Traditional fluid simulation methods exhibit a high level of simulation fidelity [25, 2, 30, 32, 12], but they often entail complex computations and significant computational expenses, thereby constraining their utility in situations that require real-time predictions. In contrast, neural networks have emerged as a highly promising approach to capturing the intricacies of natural phenomena from data without incurring enormous computational costs [14, 26, 17].
A crucial consideration in employing neural networks for physics-based learning is selecting the appropriate data representation. Point clouds are widely employed in computer vision and robotics applications to represent the 3D environment[20, 21, 28, 4]. PointNet[20], which directly applies deep learning models to point cloud data, is an excellent feature extractor. However, for complex and dynamic 3D fluid simulations, a more sophisticated approach is required for accurate computation. A common approach is to utilize graph structures [13, 1, 24, 18] to create a dynamic graph connecting nearby particles, facilitating the modeling of fluid behavior. This is achieved by updating nodes and edges to simulate ’message passing’ between SPH particles. Graph-based methods are strongly tied to particle discretization, but a wide range of physical phenomena, such as fluid mechanics, are governed by continuous partial differential equations (PDE) instead of discrete graph structures.
Most related to our approach is [27, 19]. [27] proposed a continuous convolution kernel (CConv) designed for establishing connections between particles and their neighbors. CConv exhibits strong learning capabilities for fluid particles and can broadly capture the dynamics of fluid motion. However, due to the absence of explicit physical constraints, its predictions may deviate from physical laws. Therefore, [19] has ingeniously designed an anti-symmetric convolution kernel building upon the continuous convolution kernel from CConv to introduce a robust constraint, allowing the network to approximate the conservation of momentum in physics during training. However, this strong constraint significantly complicates the problem for the network, leading to a reduction in its learning capacity. Single-path networks always confront this dilemma, while our dual-pipeline network excels at balancing these two distinct learning approaches, enabling our network to achieve a better equilibrium between global fluid control and adherence to physical law constraints.
3 Method
Figure 2: The architecture of our network. It consists of two pathways: the Main Path and the Constraint-Guided Path. These two pathways share a similar structure, with the Main Path utilizing CConv as its convolutional kernel (the orange rectangle), and the Constraint-Guided Path employing ASCC as its kernel (the blue diamond). At each layer, both pathways pass through a module for feature fusion before feeding the fused features into the next layer.
Algorithm 1 Simulation Process
  $\bm{Input:}x^{t}=\left\{x_{1},x_{2},...x_{N}\right\},v^{t}=\left\{v_{1},v_{2},...v_{N}\right\}$
  for all particles i do
     apply external forces $v_{i}^{*}\Leftarrow v_{i}+\bigtriangleup t\frac{F_{ext}}{m}$
     update position $x_{i}^{*}\Leftarrow x_{i}+\Delta tv_{i}^{*}$
  end for
  $x^{t*}=\left\{x_{1}^{*},x_{2}^{*},...x_{N}^{*}\right\}$
  $v^{t*}=\left\{v_{1}^{*},v_{2}^{*},...v_{N}^{*}\right\}$
  predict position $\Delta x^{t+1}=M_{\theta}\left(x^{t*},v^{t*}\right)$
  update position $x^{t+1}=x^{t*}+\Delta x^{t+1}$
  update velocity $v^{t+1}=\frac{x^{t+1}-x^{t}}{\Delta t}$
  $\bm{Output:}x^{t+1},v^{t+1}$
Our network follows the position-based fluids (PBF) [16] scheme. Algorithm 1 outlines the general procedure of this approach, where $M_{\theta}$ represents our neural network model with trainable parameters $\theta$, and $\Delta t$ represents the time interval between two consecutive time steps. Given a set of particles, denoted as $p_{N}$, where $N$ represents the particle number, the algorithm proceeds to analyze their positions and velocities. It takes $x^{t}$ and $v^{t}$ as inputs, where $x^{t}$ represents the positions of the particles at time step $t$ and $v^{t}$ represents their velocities. Our central learning goal is to approximate the underlying physical dynamics to predict the state of particles at the next time step $t+1$.
3.1 Two Types of Convolution Kernels
We introduce two types of convolution kernels, denoted as CConv[27] and ASCC[19]. Given a point cloud $p_{N}$ containing $N$ points labeled as $i=1,...,N$, with corresponding values $f_{i}$ at positions $x_{i}$, the convolution of CConv at position $x\in p_{n}$ can be defined as follows:
$CConv_{g}$ $=\left(f*g\right)\left(x\right)$ (1)
$=\sum_{i\in\mathcal{N}\left(x,R\right)}a\left(x_{i},x\right)f_{i}g\left(\Lambda\left(x_{i}-x\right)\right).$
Where $\mathcal{N}\left(x,R\right)$ represents the set of points located within a radius R around the position x. $g$ is the filter function, utilizing a mapping $\Lambda(r)$ that transforms a unit ball into a unit cube, described by [7, 27]. $a$ is a window function applied for density normalization specific to the points $x_{i}$ and $x$ as in [10]:
$a\left(x_{i},x\right)=\left\{\begin{array}[]{lc}\left(1-\frac{\left\|x_{i}-x_{2}^{2}\right\|}{R^{2}}\right)^{3}&\left\|x_{i}-x\right\|_{2}<R\\ 0&else.\\ \end{array}\right.$ (2)
Figure 3: The anti-symmetric kernel is obtained by negating and mirroring the trainable variables by the center point.
ASCC can be regarded as a specialized variant of CConv, where it enforces a strong constraint through the antisymmetric kernel design, allowing it to learn fluid features while adhering to the conservation of momentum. This is achieved by halving the learnable kernel parameters along a chosen axis and determining the second half through reflection about the center of the kernel, shown in Figure 3. Notably, the mirrored values are negated. ASCC can be defined as:
$ASCC_{g_{s}}$ $=\left(f*g_{s}\right)\left(x\right)$ (3)
$=\sum_{i\in\mathcal{N}\left(x,R\right)}a\left(x_{i},x\right)(f+f_{i})g_{s}\left(\Lambda\left(x_{i}-x\right)\right).$
Where $g_{s}$ represents the anti-symmetric convolution kernel and $f$ represents the corresponding value at the position $x$, with the remaining symbols defined similarly to the definitions in Equation 1.
3.2 Dual-pipeline Architectrue
Our network architecture is illustrated in Figure 2. Instead of manually determining the proportion of importance for the two pathways, we employ an innovative Attention-based Feature Fusion Module, allowing the network to autonomously learn the optimal connection between the two pathways. Instead of manually assigning specific weights to each pathway, we utilize an innovative Attention-based Feature Fusion Module, shown as Figure 3(a), enabling the network to autonomously learn the optimal fusion between the two pathways.
More precisely, our network comprises a main pathway and a constraint-guided pathway. Both pathways share a similar structure, each comprising a Type-aware Input Module and a sequence of continuous convolutions with an effective depth of five. The main pathway, which utilizes CConv as its kernel, is responsible for establishing a global context and stabilizing the fluid. The Constraint-guided pathway employs Anti-symmetric Convolution as its kernel, which enforces a strong constraint to ensure momentum conservation while learning fluid motion patterns. Before feeding the input into the convolutional sequence, we combine the result from the input module and a fully-connected layers by addition. In the second and fourth layers of the convolution sequence, we leverage the Feature Fusion Module to implement a residual connection concept [9]. We found that these designs improve accuracy and robustness.
(a) Attention-based Feature Fusion
(b) Particle Selector
Figure 4: The schematic of the Attention-based Feature Fusion Module. Features from both pathways pass through two layers of CConv and are then concatenated. The concatenated features are subsequently input into a Particle Selector to determine the respective weights for fusion. The original features from both pathways are then combined with weighted fusion.
Our network seamlessly fuses these two pathways through an Attention-based Feature Fusion Module to achieve a harmonious balance between the two pathways and unlock their maximum potential. In each layer of the convolutional sequence, assuming the features from the Main Path as $f_{main}$ and the features from the Constraint-guided Path as $f_{cons}$, the fusion of these two features in the Feature Fusion Module can be represented as follows:
$f_{fusion}$ $=S(\phi(f_{main})\oplus\phi(f_{cons}))\otimes f_{cons}$ (4)
$+(1-S(\phi(f_{main})\oplus\phi(f_{cons})))\otimes f_{main}.$
Where $\phi$ represents the $CConv$ function(Equation 1). $S$ represents the Attention-based Particle Selector and its architecture is illustrated in the 3(b). With the concatenated features from the two pathways as input, the selector outputs a selection weight between 0 and 1, which is used to assess the respective significance of features from both pathways for fusion, aiming to achieve an optimal feature fusion. Unlike previous approaches, this kind of Attention-based fusion method allows our network to learn the optimal fusion between the two pathways autonomously, endowing our network with strong learning capabilities while ensuring compliance with the law of momentum conservation.
3.3 Type-aware Input Module (TaIM)
Figure 5: The schematic of the Type-aware Input Module in the Constraint-guided Path, with ASCC as the kernel. In contrast, the Main Path’s Type-aware Input Module replaces the ASCC kernel with CConv. The Particle Selector is the same as Figure 3(b).
We propose an innovative approach to facilitate better differentiation between fluid and solid particles in the input, promoting a better understanding of fluid-solid coupling. Consequently, this enables the network to more accurately and reliably compute their interactions during collisions. Figure 5 provides an example of the Type-aware Input Module architecture in the Constraint-guided Path, with ASCC as the kernel. In contrast, the Type-aware Input Module in the Main Path replaces the ASCC kernel with the CConv kernel. Unlike the Attention-based Feature Fusion Module between two pathways, the input features here represent two entirely different types, posing a greater challenge for our network to recognize their respective diverse properties. The quality of the initial fusion can significantly influence the final fusion weights. Since it is still a feature fusion problem, an intuitive approach is to employ an additional attention module for fusing these input features, which can be represented as:
$f_{coup}^{1}$ $=S(\psi(f_{fluid})\oplus\psi(f_{solid}))\otimes f_{fluid}$ (5)
$+(1-S(\psi(f_{fluid})\oplus\psi(f_{solid})))\otimes f_{solid}.$
$f_{coup}^{2}=S(f_{coup}^{1})\otimes f_{fluid}+(1-S(f_{coup}^{1})\otimes f_{solid}.$ (6)
Where $\psi$ represents the $ASCC$ function (Equation 3). $f_{coup}^{1}$ represents the output of the first attention module, while $f_{coup}^{2}$ serves as the module’s final output.
3.4 Training Procedure
Consistent with prior research [27, 13, 24], we employ a loss function that calculates the mean absolute error in position values between the predicted and ground truth weighted by the neighbor count as our loss function:
$\mathcal{L}^{n+1}=\sum_{i=1}^{N}\varphi_{i}\left\|x_{i}^{t+1}-\hat{x}_{i}^{t+1}\right\|_{2}^{\gamma}.$ (7)
Where $x_{i}^{t+1}$ represents the predicted position from the Algorithm 1 at time step $t+1$, and $\hat{x}_{i}^{t+1}$ represents the ground-truth position. $\varphi_{i}$ is the weight of neighbor count to emphasize the loss for particles with fewer neighbors:
$\varphi_{i}=exp(-\frac{1}{c}\left|\mathcal{N}(x_{i}^{t*})\right|).$ (8)
$\mathcal{N}(x_{i}^{t*})$ represents the neighbor count of the particle $i$ at time step $t$ after experiencing external force. We set $c=40$, in accordance with the average number of neighboring particles and $\gamma=0.5$ following [27], which makes the loss function more sensitive to small particle motions. For temporal stability considerations, we conduct training to predict particle positions for two future time steps, $t+1$ and $t+2$. This extension contributes to an overall enhancement in simulation quality. As a result, the combined loss function can be represented as follows:
$\mathcal{L}=\mathcal{L}^{t+1}+\mathcal{L}^{t+2}.$ (9)
4 Experiments
Method Average pos err(mm) Average distance to Wasserstein Inference
n+1 n+2 closest point $d^{n}$(mm) distance time(ms)
Liquid(6k) DPI-Nets [13] 26.19 51.77 unstable - 305.55
KPConv [26] 1.65 4.54 unstable - 57.89
CConv [27] 0.56 1.51 29.50 0.26 16.47
DMCF [19] 0.65 1.89 32.89 0.21 94.86
Ours 0.43 1.16 28.32 0.17 48.01
DamBreak DPI-Nets [13] 12.73 25.38 22.07 - 202.56
KPConv [26] 2.49 7.05 unstable - 47.96
CConv [27] 0.57 1.38 20.63 0.21 12.01
DMCF [19] 0.64 1.83 24.56 0.20 70.49
Ours 0.42 1.02 20.90 0.18 35.57
Table 1: Quantitative Evaluation. We evaluated the networks according to the evaluation metrics described in 4.2, and also compared the inference time required per frame for each network. It’s worth noting that some methods become unstable after a few frames.
Figure 6: Qualitative Evaluation. In the results of CConv, the particles exhibit a clear overfitting to gravity. DMCF can effectively capture the physics of falling, but its ability to learn stable overall fluid dynamics is somewhat lacking, leading to occasional distortions at the edges. Our network better adheres to physical laws and reduces simulation errors, bringing it closer to the ground truth.
Our experiments aimed to showcase the prediction accuracy, probability distribution divergence, and inference time of our approach compared to various baseline methods. What’s more, we show the great generalization and broad applicability of our network with adequate visual analysis.
4.1 Datasets
We conducted experiments on three distinct 3D fluid datasets. Two of these datasets were derived from the Liquid3D dataset presented by [27]. The first dataset, denoted as ’Liquid(6k),’ represents a simpler scenario with a fixed particle count of 6000. The second dataset, denoted as ’Liquid(complex),’ encompasses a variety of complex scenarios featuring different initial shapes of fluid particles and multiple complex boxes, with particle counts ranging from thousands to tens of thousands. The Liquid3D dataset was generated using DFSPH [2], which emphasizes high-fidelity simulations and operates at a 50 Hz time resolution. It comprises 200 training scenes and 20 test scenes. The third dataset, ’DamBreak,’ was from [13] and consists of 2000 training scenes and 300 testing scenes. These scenes simulate the behavior of a randomly positioned fluid block in a static box.
4.2 Evaluation metrics
To evaluate the prediction accuracy of each method, we calculate the average positional error by comparing the predicted positions to the ground truth. We start the assessment by using every $5^{th}$ frame for initialization and compute the deviations from the ground truth between two consecutive frames, denoted as $t+1$ and $t+2$. Furthermore, we measure long-term similarity by reporting the average distance from the ground-truth particles to the closest particle in the entire sequence. This distance for frame $n$ is computed as follows:
$d^{n}=\frac{1}{N}\sum_{i=1}^{N}\min_{x^{n}\in X^{n}}\left\|\hat{x}_{i}^{n}-x^{n}\right\|_{2}.$ (10)
Where $X^{n}$ represents the set of predicted particle positions for frame $n$, $\hat{x}_{i}^{n}$ corresponds to the ground-truth position for particle $i$, and $N$ denotes the total number of particles. We quantify the divergence between the prediction and ground truth probability distributions by calculating the Wasserstein Distance [23] between them. The Wasserstein Distance, also known as the Earth Mover’s Distance (EMD), is a measure of the transportation cost required to transform one probability distribution into another, which can be used to measure the similarity between two distributions.
4.3 Comparative Experiments
Method Average pos err(mm) Average distance to Wasserstein
t+1 t+2 closest point $d^{n}$(mm) distance
Main Path (no TaIM) 0.66 1.86 30.83 0.26
Main Path 0.58 1.61 30.67 0.21
Constraint-guided Path (no TaIM) 0.75 2.18 36.94 0.30
Constraint-guided Path 0.67 1.86 32.20 0.25
Ours 0.52 1.45 29.42 0.19
Table 2: Ablation Study. Our ablation experiments were conducted on the ’Liquid (complex)’ dataset, testing the following configurations: a single Main Path without TaIM (Type-aware Input Module), a single Main Path with TaIM, a single Constraint-guided Path without TaIM, a single Constraint-guided Path with TaIM, and our final Dual-pipeline network.
Method Average pos err(mm) Average distance to Wasserstein
t+1 t+2 closest point $d^{n}$(mm) distance
CConv 0.66 1.86 30.83 0.26
CConv-TaIM 0.58 1.61 30.67 0.22
DMCF 0.73 2.08 47.89 0.21
DMCF-TaIM 0.69 2.03 46.58 0.21
Ours (no TaIM) 0.57 1.59 30.04 0.20
Ours 0.52 1.45 29.42 0.19
Table 3: The plug-and-play usability of TaIM. We incorporated TaIM as a preprocessing step before input in CConv, DMCF, and our network, and compared the performance differences with and without TaIM. We can easily integrate TaIM into networks with similar convolution kernels, resulting in a significant performance improvement. This demonstrates that the concept behind TaIM is beneficial for fluid-solid coupling.
Figure 7: Generalization to fluid initial shapes: the Stanford Bunny fluid free-falls into a box with a circular depression at the bottom, creating splashes. See the supplementary video.
Figure 8: Generalization to complex terrains and varying fluid particle scenarios: fluid emitted from a particle emitter flows from a higher elevation in a complex valley, eventually filling the depressions in a plain. See the supplementary video.
Quantitative Evaluation
Table 1 presents comparative experiments conducted on the Liquid(6k) and DamBreak datasets. We assess the average error between the ground-truth particle positions and the two predicted future frames in the test set. Additionally, we calculate the average distance from the ground truth to the prediction throughout the entire sequence. The Wasserstein Distance is employed to gauge the similarity between the predicted particle distribution and the ground truth distribution.
We compare our method to DPI-Nets [13], two kinds of continuous convolution network [26, 27] and DMCF [19]. It is worth noting that DMCF exhibits an increase in mean squared error (MSE) compared to CConv [27]. This is because the anti-symmetric convolution layer introduces a strong constraint, which makes the problem considerably more challenging for the network to solve. Merely appending the ASCC at the end made it more complex to fine-tune the network to its current state. As a trade-off, the network exhibits enhanced generalization. What’s more, the Wasserstein Distance to the ground truth has also shown a decrease, indicating a distribution that more closely approximates the physically accurate ground truth. Our network outperforms in both prediction accuracy and probability distribution divergence compared to previous models. This achievement is attributed to the design of our Dual-pipeline architecture, which attains equilibrium between global fluid control and adherence to physical law constraints.
Qualitative Evaluation
Figure 6 presents a visual comparison of our network with CConv and DMCF on the Liquid3D dataset. To better analyze the underlying physical principles during the dynamic simulation process, we selected five frames from the free-fall to collision with the container’s bottom for analysis. CConv starts to deform its structure from the second frame onwards, which significantly deviates from the expected physical laws of free fall. This is attributed to CConv’s lack of emphasis on the law of momentum conservation, allowing it to overfit gravity and negate it. This reduces the problem to a local per-particle problem. In contrast, DMCF exhibits an improved adherence to the laws of physics during the free fall phase, thanks to the inclusion of ASCC. However, DMCF’s approach of simply adding ASCC as the final output layer compromises the overall stability of the fluid. This instability becomes evident from the second frame, where the edges of the cube exhibit noticeable distortion. On the other hand, our network excels in simulating the fluid by providing a stable global context while maintaining a strong grasp of physical principles throughout the simulation process, bringing it closer to the ground truth.
4.4 Ablation Study
We conduct an ablation study to assess the design choices of our network, shown in Table 2. This study employs the same evaluation metrics as the comparative experiments. We systematically analyze the contributions of each pathway independently and evaluate the enhancements provided by the Type-aware Input Module to both pathways. In addition, we also confirm the plug-and-play capability of the Type-aware Input Module. We conducted separate tests by adding this module to CConv, DMCF, and our network, as they all utilize similar convolution kernels. As illustrated in Table 3, this module proves to be instrumental in enhancing fluid simulation accuracy, indicating its pivotal role in achieving fluid-solid coupling.
4.5 Generalization
We validate the generalization capability of our network to unseen scenarios from two perspectives. Firstly, Figure 7 presents the generalization to irregular initial shapes (Stanford Bunny), which is completely unseen and full of challenges, as the initial shapes in the Liquid3D dataset are all regular. Secondly, we simulated a scenario where fluid, emitted from a particle emitter, flows from a high point in a complex valley terrain down to fill a depression in a plain, demonstrating the network’s generalization ability regarding scene geometry and the number of particles, as shown in Figure 8. The results demonstrate the strong generalization capabilities of our method in handling complex liquid shapes and terrains. This is attributed to our Dual-pipeline architecture, which effectively learns fluid dynamics while tightly adhering to physical laws.
5 Conclusion
We have designed an Attention-based Dual-pipeline network architecture, which provides stable global fluid control and ensures adherence to physical law constraints in fluid simulation. The Dual-pipeline feature fusion architecture also holds the potential to inspire work in other domains. Additionally, we propose a Type-aware Input Module, which plays a pivotal role in facilitating the fluid-solid particle coupling. Our network represents a step forward in utilizing neural networks to simulate physically compliant fluids, enhancing the potential for real-time simulation generalization across diverse fluid scenarios, particularly in scenarios such as interactive gaming environments, and virtual reality simulations. There remains untapped potential in enabling neural networks to learn real-world physics. We will make our code publicly available to contribute to the ongoing research.